Nominal income, price level and purchasing power Suppose Jane gets a raise in salary of 10%...

  1. Nominal income, price level and purchasing power

Suppose Jane gets a raise in salary of 10% and she finds that her cost of living (the price level) has increased by 15%, what happens to her purchasing power?
Answer #1

We can do this with the help of a example.

Suppose if Jane initially earns $200

After 10% his income becomes $220

Suppose the price of a shirt is $100

After 15 % increase, the new price becomes $115

So we can see initially Jane could buy (200/100)=2 shirts and he can buy only (220/115)=1.913 shirts

So now , her purchasing power decreases by ((2-1.913)/2)*100 = 4.545%

We can say that Jane’s purchasing power decreases and it decreases by 4.545%